# Connecting essential triangulations II: via 2-3 moves only

### v1

@@ Abstract.
In previous work we showed that for a manifold $M$, whose universal cover has infinitely many boundary components, the set of essential ideal triangulations of $M$ is connected via 2-3, 3-2, 0-2, and 2-0 moves. Here we show that this set is also connected via 2-3 and 3-2 moves alone, if we ignore those triangulations for which no 2-3 move results in an essential triangulation. If we also allow V-moves and their inverses then the full set of essential ideal triangulations of $M$ is once again connected. These results also hold if we replace essential triangulations with $L$--essential triangulations.
@@ 1. Introduction
1.1. Paths of triangulations
The various notions of equivalence of triangulations have been refined many times, notably by Pachner [6], who showed that bistellar moves connect any two triangulations of a given manifold. For a three-dimensional manifold $M$, these are the 2-3, 3-2, 1-4, and 4-1 moves. If we do not want to change the number of vertices, the latter two moves are almost always unnecessary, by the work of Matveev [5, Theorem 1.2.5], Piergallini [7, Theorem 1.2], and Amendola [1, Theorem 2.1]. To state their results, let $\mathbb{T}(M)$ be those one-vertex triangulations of $M$ (if $M$ is closed) or ideal triangulations (if $M$ has boundary). Let $\mathbb{T}_{2}(M)$ be those triangulations of $\mathbb{T}(M)$ having at least two tetrahedra.
Theorem 1.1.
@@ 1.1. Paths of triangulations
An essential ideal triangulation has the property that none of its edges are homotopic, relative to their endpoints, into $\partial M$. (In this paper we will only consider ideal triangulations of three-manifolds with boundary.) Essentiality is required in many applications. For example, an ideal triangulation of a cusped hyperbolic three-manifold is essential if and only if it admits a solution to Thurston’s gluing equations.
In the first paper in this series we prove a connectivity result similar to Theorem 1.1 but where the initial and terminal triangulations, as well as all intermediate triangulations, are essential. Also in that paper we generalise to $L$--essential triangulations. These involve a choice of labelling $L$ of the boundary components $\Delta_{M}$ of the universal cover of the manifold $M$. See Definition 2.6 for the details. Let $\mathbb{T}(M,L)$ be those (necessarily ideal) triangulations in $\mathbb{T}(M)$ which are $L$--essential. The full result [3, Theorem 6.1] is as follows.
@@
This result makes use of 0-2 and 2-0 moves to connect triangulations together. These are again local moves on the triangulation but are not bistellar moves. See Section 2.3.1. Let $\mathbb{T}^{\circ}(M,L)$ be the triangulations in $\mathbb{T}(M,L)$ which admit some 2-3 move preserving $L$--essentiality. The main goal of this paper is to prove the following.
@@
Thus we improve on Theorem 1.2 by removing the need for 0-2 and 2-0 moves, at the cost of being unable to connect to certain isolated $L$--essential triangulations. These isolated triangulations include those excluded in Theorem 1.1: that is, triangulations with only one tetrahedron. However, more subtle isolations occur in the context of $L$--essential triangulations. See Example 2.10.
On the other hand, each $L$--essential triangulation can be connected to $\mathbb{T}^{\circ}(M,L)$ by a single V-move, a more ‘‘local’’ kind of 0-2 move. See Definition 3.8 and Lemma 3.10. Combining Theorem 3.5 and Lemma 3.10, we obtain the following improvement of Theorem 1.2.
@@
The second and third authors will use Corollary 1.3 in their proof that the figure-eight knot complement, and certain other manifolds, have unique veering triangulations. In the proof, we carry the data of a circular order through a sequence of 2-3 moves, V-moves, and their inverses. Surprisingly, in this setting the 2-0 move is problematic. In particular, Theorem 1.2 does not suffice for our future work.
@@ 2. Background
Here we essentially follow the notation and definitions of the first paper in this series [3, Section 2].
@@ 2.2. Labellings and $L$--essentiality
See [3, Section 2.9] for examples of labellings. The simplest labelling is the identity map on $\Delta_{M}$. As noted in [3, Remark 2.18], with this labelling $L$--essential triangulations are essential triangulations in the sense of [2, Definitions 3.2 and 3.5] and [4, page 336].
@@ 2.3. Moves on foams
The three-dimensional bistellar moves [6] are the 1-4, 2-3, 3-2, and 4-1 moves. These can be performed equally well on triangulations or their dual foams. Of these we only consider the 2-3 and 3-2 moves, applied to foams. The former is called the T move by Matveev [5, page 14]. It can be performed along any edge of $\mathcal{F}$ that is not an edge loop. See Figure 2.2. The 3-2 move can be performed on any triangular face whose closure is embedded in $\mathcal{F}$.
@@ 2.3.1. The 0-2 move
The 0-2 move, defined by Figure 2.3, is called the ambient lune move by Matveev [5, page 17]. It is applied along an arc $\delta$ properly embedded in, and avoiding the vertices of, a face of $\mathcal{F}$. The 0-2 move creates two new vertices and a new bigon face. We denote the result by $\mathcal{F}[\delta]$.
@@ 2.4. Edge loops cause issues
Here we show that one cannot remove 0-2 and 2-0 moves from the statement of Theorem 1.2 without adding some other hypothesis.
@@
Figure 2.4a shows a foam in a solid torus. Mirroring this foam across the boundary torus produces a foam in $S^{2}\times S^{1}$ with two vertices, four edges, three faces, and one complementary region. Let $L$ be the identity labelling (as described in Section 2.2). All faces are essential (in the universal cover the regions to either side of each face are distinct). There are no triangular faces so no 3-2 move is possible. Two of the edges are edge loops so a 2-3 move cannot be applied along them. The other two edges bound a bigon face, which implies that a 2-3 move applied along them introduces an inessential face.
Applying various 0-2 moves produces other foams with all faces essential. Depending on which 0-2 move we apply, the resulting foam may or may not have a 2-3 move keeping the foam essential. For example, suppose that we perform a 0-2 move across the bigon face. Then in the resulting foam no 2-3 or 3-2 move is possible while maintaining essentiality.
Therefore there are at least some essential triangulations that are isolated if we can only use 2-3 and 3-2 moves. ∎
This is a similar situation to Matveev’s, Piergallini’s, and Amendola’s original result (Theorem 1.1) in which triangulations with a single tetrahedron are isolated.
3. Implementing 0-2 moves
@@
If $L$ has infinite image then $\mathbb{T}(M,L)$ is non-empty by [3, Theorem 3.1]. Lemma 3.10 then implies that $\mathbb{T}^{\circ}(M,L)$ is non-empty. ∎
@@ 3.2. V-moves
We use some of the same tools as Matveev, beginning with the V-move [5, Definition 1.2.6].
@@
Suppose that $\mathcal{F}$ is a foam. Suppose that $v$ is a vertex of $\mathcal{F}$. Let $\mathcal{N}$ be a small regular neighbourhood of $v$. Suppose that $\delta_{+}$ and $\delta_{-}$ are a pair of opposite edges of $\partial(\mathcal{F}\cap\mathcal{N})$. Then the V-move on $v$ (in the direction of $\delta_{+}$ and $\delta_{-}$) is the 0-2 move along $\delta_{+}$. (Equivalently, it is also the 0-2 move along $\delta_{-}$). ∎
@@ 3.3. Avoiding cyclic edges
By Lemma 3.6, there is an edge $e$ of ${\widetilde{\mathcal{F}}}$ that is incident to complementary regions with five distinct labels. Since $\mathcal{F}$ contains no $L$–flippable edges, $\phi_{M}(e)$ must be cyclic. Let $\mathcal{N}$ be a small regular neighbourhood of $v$, the unique endpoint of $\phi_{M}(e)$. Let $\delta_{+}$ be as given in Definition 3.8. We obtain $\mathcal{F}^{\prime}$ by performing a 0-2 move along $\delta_{+}$. By Lemma 3.9, the foam $\mathcal{F}^{\prime}$ is $L$–essential. Also, the descendant of $\phi_{M}(e)$ not contained in $\mathcal{N}$ is $L$–flippable. ∎
Also using notation as in Definition 3.8, we have the following.
Lemma 3.11.
Suppose that $e$ is a cyclic edge of $\mathcal{F}$ with both endpoints at $v$. Suppose that exactly one endpoint of $\delta_{+}$ lies on $e$. Then the V-move at $v$ obtained by applying a 0-2 move along $\delta_{+}$ gives a foam $\mathcal{F}^{\prime}$ with fewer cyclic edges than $\mathcal{F}$.
Proof.
The cyclic edge $e$ is destroyed by the vertices added by the 0-2 move. By Lemma 3.9, no new cyclic edges are created. ∎
@@
The proof of Proposition 3.14 is quite difficult and takes up the remainder of the paper.
@@ 4. Following Matveev
We will give a series of constructions that prove Proposition 3.14 under increasingly general circumstances. Following [5, Proposition 1.2.8], we begin with the construction of a V-move in Lemma 4.1, where the arc $\delta$ that the 0-2 move is to be applied along cuts off a single vertex $v$ of its model face.
@@
Figure 4.1. A V-move can be implemented by three 2-3 moves followed by a 3-2 move. Adapted from [5, Figure 1.15].
@@ 4.1. Moving snakelets to realise some 0-2 moves
Let $A=E_{0}$ and $B=E_{2}$. See Figure 4.4a. Let $N$ and $S$ be the regions of ${\widetilde{M}}-{\widetilde{\mathcal{F}}}$ incident to the interior of $f$. The labels on $A$, $B$, $E$, $N$, and $S$ are all distinct because $\mathcal{F}[\delta]$ is $L$–essential. Because $L(A)\neq L(B)$ and $e_{1}$ is not cyclic, we have that $e_{1}$ is $L$–flippable. By Lemma 4.1 we may perform a V-move at the vertex $v_{0}$ where $e_{0}$ meets $e_{1}$. See Figure 4.4b. This produces the red snakelet. Its head is on $e_{1}$ and its tail is on $e_{0}$. (Here we think of $e_{1}$ and $e_{0}$ as subsets of $M$, since after the V-move, they are no longer edges of the foam.) Since $\phi_{M}(e_{1})$ is distinct from $\phi_{M}(e_{0})$, we may isotope the head of the red snakelet to the other end of $e_{1}$. Applying Lemma 4.3, we move the head of the red snakelet onto $e_{2}$. At this stage we are done unless we are in the case that $\phi_{M}(e_{0})=\phi_{M}(e_{2})$, and the head and the tail of the red snakelet are in the wrong order along the edge $e_{2}$. (Matveev swaps the ends of a snakelet in [5, Figure 1.19]; maintaining $L$–essentiality requires additional work as follows.)

### v2

@@ Abstract.
In previous work we showed that for a manifold $M$, whose universal cover has infinitely many boundary components, the set of essential ideal triangulations of $M$ is connected via 2-3, 3-2, 0-2, and 2-0 moves. Here we show that this set is also connected via 2-3 and 3-2 moves alone, if we ignore those triangulations for which no 2-3 move preserves essentiality. If we also allow V-moves and their inverses then the full set of essential ideal triangulations of $M$ is once again connected. These results also hold if we replace essential triangulations with $L$--essential triangulations.
@@ 1. Introduction
1.1. Graphs of triangulations
Combinatorial moves on triangulations have been studied extensively. A particularly elegant formulation was given by Pachner [9], who showed that bistellar moves connect any two triangulations of a given manifold. For a three-dimensional manifold $M$, these are the 2-3, 3-2, 1-4, and 4-1 moves. If we do not want to change the number of vertices, we may restrict ourselves to the former two moves by the work of Matveev [7, Theorem 1.2.5], Piergallini [10, Theorem 1.2], and Amendola [1, Theorem 2.1]. To state their results, let $\mathbb{T}(M)$ be the set of one-vertex triangulations (if $M$ is closed) or the set of ideal triangulations (if $M$ has boundary). Let $\mathbb{T}_{2}(M)$ be those triangulations of $\mathbb{T}(M)$ having at least two tetrahedra.
Figure 1.1. Part of the graph having $\mathbb{T}_{2}(S^{2}\times S^{1})$ as vertices (triangulations arranged by number of tetrahedra) and 2-3 moves as edges. This and the following graphs were generated using Regina [2].
Theorem 1.1 (Matveev–Piergallini–Amendola).
@@ 1.1. Graphs of triangulations
We illustrate this for $S^{2}\times S^{1}$ in Figure 1.1. An essential ideal triangulation has the property that none of its edges are homotopic, relative to their endpoints, into $\partial M$. (When $M$ is closed, we require instead that none of its edges are null-homotopic. Equivalently, by removing regular open neighbourhoods of the vertices, we may convert material triangulations into ideal ones.) Essentiality is required in many applications. For example, an ideal triangulation of a cusped hyperbolic three-manifold is essential if and only if it admits a solution to Thurston’s gluing equations [12, Theorem 1]. (Here we do not assume that all tetrahedra are positively oriented.) Thus the 1-loop invariant of Dimofte and Garoufalidis [3, Definition 1.2] is only defined on essential triangulations, as is the Bloch invariant [8, Definition 2.5]. Moreover, a triangulation admitting a strict angle structure or a taut angle structure is necessarily essential [4, Theorem 6.1]. Unfortunately, the set of essential triangulations need not be connected via 2-3 and 3-2 moves. This is the case for $S^{2}\times S^{1}$, as shown in Figure 1.2.
Figure 1.2. The subgraph of Figure 1.1 spanned by essential triangulations. For a discussion of the isolated essential triangulation with two tetrahedra, see Example 2.10.
In the prequel [5] to this paper we prove a connectivity result similar to Theorem 1.1 but where the initial and terminal triangulations, as well as all intermediate triangulations, are essential. To do this, we use 0-2 and 2-0 moves; see Section 2.3.1. The resulting connected graph (again for $S^{2}\times S^{1}$) is shown in Figure 1.3.
Figure 1.3. The graph of Figure 1.2 with additional edges for 0-2 moves.
Also in that paper we generalise to $L$--essential triangulations. These involve a choice of labelling $L$ of the boundary components $\Delta_{M}$ of the universal cover of the manifold $M$. See Definition 2.6 for the details. Let $\mathbb{T}(M,L)$ be those (necessarily ideal) triangulations in $\mathbb{T}(M)$ which are $L$--essential. The full result [5, Theorem 6.1] is as follows.
@@
Let $\mathbb{T}^{\circ}(M,L)$ be the triangulations in $\mathbb{T}(M,L)$ which admit some 2-3 or 3-2 move preserving $L$--essentiality. The main goal of this paper is to prove the following.
@@
This implies that in the graph illustrated in Figure 1.2 there is only one component with edges. We say that a triangulation is isolated if it lies in $\mathbb{T}(M,L)-\mathbb{T}^{\circ}(M,L)$.
As it happens, every isolated triangulation can be connected to $\mathbb{T}^{\circ}(M,L)$ by a single V-move, a more ‘‘local’’ 0-2 move. See Definition 3.8 and Lemma 3.14. Using (selected) V-moves instead of the more general 0-2 moves yields the connected graph shown in Figure 1.4.
Figure 1.4. Essential one-vertex triangulations of $S^{2}\times S^{1}$, as connected by 2-3 and (selected) V-moves.
Combining Theorem 3.5 and Lemma 3.14, we obtain the following improvement of Theorem 1.2.
@@
In [11], the second and third authors will use Corollary 1.3 to classify the veering triangulations of hyperbolic once-punctured torus bundles and various other manifolds. To do this, we carry a ‘‘winding’’ and a compatible circular ordering through a sequence of essential triangulations, connected by 2-3 moves, V-moves, and their inverses. Surprisingly, some windings cannot be carried through certain 2-0 moves. Thus Theorem 1.2 does not suffice for this application.
@@ 2. Background
Here we essentially follow the notation and definitions of [5, Section 2].
@@ 2.2. Labellings and $L$--essentiality
See [5, Section 2.9] for examples of labellings. The simplest labelling is the identity map on $\Delta_{M}$. As noted in [5, Remark 2.18], with this labelling $L$--essential triangulations are essential triangulations in the sense of [4, Definitions 3.2 and 3.5] and [6, page 336].
@@ 2.3. Moves on foams
The three-dimensional bistellar moves [9] are the 1-4, 2-3, 3-2, and 4-1 moves. These can be performed equally well on triangulations or their dual foams. Of these we only consider the 2-3 and 3-2 moves, applied to foams. The former is called the T move by Matveev [7, page 14]. It can be performed along any edge of $\mathcal{F}$ that is not an edge loop. See Figure 2.2. The 3-2 move can be performed on any triangular face whose closure is embedded in $\mathcal{F}$.
@@ 2.3.1. The 0-2 move
The 0-2 move, defined by Figure 2.3, is called the ambient lune move by Matveev [7, page 17]. It is applied along an arc $\delta$ properly embedded in, and avoiding the vertices of, a face of $\mathcal{F}$. The 0-2 move creates two new vertices and a new bigon face. We denote the result by $\mathcal{F}[\delta]$.
@@ 2.4. Edge loops cause issues
Here we give an explicit example of an isolated essential triangulation: that is, no 2-3 or 3-2 move preserves essentiality.
@@
Figure 2.4a shows a foam in a solid torus. Mirroring this foam across the boundary torus produces a foam $\mathcal{F}$ in $S^{2}\times S^{1}$ with two vertices, four edges, three faces, and one complementary region. Let $L$ be the identity labelling (as described in Section 2.2). All faces of $\mathcal{F}$ are essential (in the universal cover the regions to either side of each face are distinct). There are no triangular faces so no 3-2 move is possible. Two of the edges of $\mathcal{F}$ are edge loops so a 2-3 move cannot be applied along them. The other two edges bound a bigon face, which implies that a 2-3 move applied along them introduces an inessential face. It follows that the triangulation $\mathcal{T}$ dual to $\mathcal{F}$ is isolated. In fact, $\mathcal{T}$ is the unique triangulation with two tetrahedra in Figure 1.2. ∎
This example is similar in spirit to the finite number of exceptions to Matveev’s, Piergallini’s, and Amendola’s original result (Theorem 1.1). There, a triangulation with a single tetrahedron is ‘‘isolated’’. However in our context, starting with $\mathcal{F}$ and repeatedly applying Remark 3.15 produces infinitely many examples.
3. $L$--flippable edges and cyclic edges
@@
Therefore, in the graph with $\mathbb{T}(M,L)$ for vertices and 2-3 moves for edges, a triangulation is isolated if and only if it has no $L$--flippable faces.
@@
If $L$ has infinite image then $\mathbb{T}(M,L)$ is non-empty by [5, Theorem 3.1]. Lemma 3.14 then implies that $\mathbb{T}^{\circ}(M,L)$ is non-empty. ∎
@@ 3.2. V-moves
We use some of the same tools as Matveev, beginning with the V-move [7, Definition 1.2.6].
@@
Suppose that $\mathcal{F}$ is a foam. Suppose that $v$ is a vertex of $\mathcal{F}$. Let $\mathcal{N}$ be a small regular neighbourhood of $v$. Suppose that $\delta_{+}$ and $\delta_{-}$ are a pair of opposite edges of $\partial(\mathcal{F}\cap\mathcal{N})$. Then the V-move at $v$ along $\delta_{+}$ is the 0-2 move along $\delta_{+}$. (Equivalently, it is also the 0-2 move along $\delta_{-}$). ∎
@@ 3.3. Avoiding cyclic edges
Suppose that $e$ is a cyclic edge of $\mathcal{F}$ with both endpoints at $v$. Then there is an arc $\delta_{+}$ as in Definition 3.8 with the additional hypothesis that it meets $e$ in exactly one point. Moreover, the V-move at $v$ along $\delta_{+}$ gives a foam $\mathcal{F}^{\prime}$ with fewer cyclic edges than $\mathcal{F}$.
@@
There are six possibilities for $\delta_{+}$ in Definition 3.8. One of these meets $e$ in two points and one of these meets $e$ in zero points. We choose one of the four remaining arcs for $\delta_{+}$. The cyclic edge $e$ is then destroyed by the vertices added by the V-move. See Figure 3.1. By Lemma 3.9, no new cyclic edges are created. ∎
@@
The proof of Proposition 3.13 is quite difficult; it begins in Section 4 and takes up the remainder of the paper.
3.4. Connecting the isolated
We can now connect isolated triangulations to $\mathbb{T}^{\circ}(M,L)$ using V-moves.
Lemma 3.14.
Proof.
By Lemma 3.6, there is an edge $e$ of ${\widetilde{\mathcal{F}}}$ that is incident to complementary regions with five distinct labels. Since $\mathcal{F}$ contains no $L$–flippable edges, $\phi_{M}(e)$ must be cyclic. Let $\mathcal{N}$ be a small regular neighbourhood of $v$, the unique endpoint of $\phi_{M}(e)$. Let $\delta_{+}$ be as given in Lemma 3.10. We obtain $\mathcal{F}^{\prime}$ by performing a 0-2 move along $\delta_{+}$. By Lemma 3.9, the foam $\mathcal{F}^{\prime}$ is $L$–essential. Also, the descendant of $\phi_{M}(e)$ not contained in $\mathcal{N}$ is $L$–flippable. ∎
Remark 3.15.
Suppose that, in the proof of Lemma 3.14, we instead choose $\delta_{+}$ to meet $e$ two or zero times. Then the resulting foam $\mathcal{F}^{\prime}$ again has no $L$–flippable edge. Thus if there is one isolated $L$–essential foam, there are infinitely many. ∎
@@ 4. Following Matveev
We will give a series of constructions that prove Proposition 3.13 under increasingly general circumstances. Following [7, Proposition 1.2.8], we begin with the construction of a V-move in Lemma 4.1, where the arc $\delta$ that the 0-2 move is to be applied along cuts off a single vertex $v$ of its model face.
@@
Figure 4.1. A V-move can be implemented by three 2-3 moves followed by a 3-2 move. Adapted from [7, Figure 1.15].
@@ 4.1. Moving snakelets to realise some 0-2 moves
Let $A=E_{0}$ and $B=E_{2}$. See Figure 4.4a. Let $N$ and $S$ be the regions of ${\widetilde{M}}-{\widetilde{\mathcal{F}}}$ incident to the interior of $f$. The labels on $A$, $B$, $E$, $N$, and $S$ are all distinct because $\mathcal{F}[\delta]$ is $L$–essential. Because $L(A)\neq L(B)$ and $e_{1}$ is not cyclic, we have that $e_{1}$ is $L$–flippable. By Lemma 4.1 we may perform a V-move at the vertex $v_{0}$ where $e_{0}$ meets $e_{1}$. See Figure 4.4b. This produces the red snakelet. Its head is on $e_{1}$ and its tail is on $e_{0}$. (Here we think of $e_{1}$ and $e_{0}$ as subsets of $M$, since after the V-move, they are no longer edges of the foam.) Since $\phi_{M}(e_{1})$ is distinct from $\phi_{M}(e_{0})$, we may isotope the head of the red snakelet to the other end of $e_{1}$. Applying Lemma 4.3, we move the head of the red snakelet onto $e_{2}$. At this stage we are done unless we are in the case that $\phi_{M}(e_{0})=\phi_{M}(e_{2})$, and the head and the tail of the red snakelet are in the wrong order along the edge $e_{2}$. (Matveev swaps the ends of a snakelet in [7, Figure 1.19]; maintaining $L$–essentiality requires additional work as follows.)